It was a fine warm day with the odd shower at Grosmont on Tuesday for Bill Dobson, Ian McCall, James Pearcy, Nigel Hall, Gordon Wells, Jon Bradley, Nigel Bill, and Ian Pearson.
As there was quite an amount of water in the pit that morning the first job was to solve a problem with the pit pump which was causing the earth leakage switch to trip. After isolating the the pump supply and then, after disconnecting the pump plug, it was found to be contaminated with water. It must have got soaked when the tender was filled the previous week. The plug and socket was cleaned and dried and, after removing some debris from the outlet bend on the pump, it shot back into action removing the water from the pit.The drainage channel could still do with another rake through however, to clear any remaining debris.
Jobs completed were:
- Righthand eccentric rods fitted secured with nuts and lock nuts.
- Lefthand forward eccentric had a 43 thou shim fitted and all the the nuts on both lefthand eccentrics were secured, including the locking nuts.
- The crosshead large cotter pin was driven out after being jammed last week.
- A lot of work was done by Gordon Wells on the new axle box R2 pump which had been giving problems. They were thought to have been resolved, but further testing the next day found it still needs more attention.
- Both sides of the tender had been varnished by Nigel Hall after he had fitted the BR logos on Monday.
- Cab side windows fitted by Jon Bradley who also made a start on cleaning and polishing the cab front windows.
- A start was made on filing the lefthand large crosshead cotter.
- A start was made on fitting eccentric rod taper pins joining the rod to the expansion links.
- Engraved brass plates were fitted to the axle box lubricators identifying oil pipe to axle box or cylinders.
Wednesday was another fine day with Bill Dobson, Ed Bolam, Nigel Hall, Chris Lawson and Ian Pearson present.
Jobs completed were:
- Righthand crosshead large cotter dressed and made to fit the crosshead slot by Nigel Hall. This took most of the day, with trial fitting and dressing throughout before finally being secured with small cotter and split pinned.
- Lefthand big end securing securing cotter made and fitted and split pinned
- Both brass front window frames polished, glass cleaned then fitted back onto the brackets in the cab by Jon Bradley, with assistance from Ed Bolam.
- Eccentric rod taper pins fitted at the expansion link end and opened out.
- Eight securing cotters fitted to the eccentric strap securing studs in front of the locking nuts, involving much pinch barring of the locomotive to get everything aligned.
- Steam reversing shaft fitted from reversing arm to oil cylinder arm.
- Some re adjustment of the new lubricator drive shaft was made with a new hole drilled and refitted.
- The tender and locomotive were re-connected at the end of the day - but only on the safety links.
The steam pipes are in the MPD running shed after machining, and a Lenz ring has been successfully trial fitted to confirm the machining is correct. They will be fitted by Barney and MPD colleagues once the Q6 is moved down to the running shed, which, for the second week running, is expected to take place on Friday ie today.
Attached are some photographs of the locomotive showing the superb finish that Nigel Hall has achieved with his paintwork, and of the interior of the cab on which Jon Bradley has concentrated - the plaques have still to be fitted to the roof and no, that is not a new chair for the driver, Jon will have to stand like the rest of the crew!